Walid has 30 books some are 4 cm thick and some are 3 cm . when arranged on a library shelf they occupy 1 m . how many books of each kind does Walid have?
Answer by mananth(16946)


Walid has 30 books some are 4 cm thick and some are 3 cm . when arranged on a library shelf they occupy 1 m . how many books of each kind does Walid have?
let there be x books of 4cm height
and y books of 3 cm height
Total books = 30
x+y = 30---------1
Height equation
4x+3y = 100----------2
Solve 1 & 2
Multiply (1) by -4
we get -4x-4y=-120
4x+3y=100 ---(2)
Add the equations
-y =-20
y=20 Therefore x = 10